the altitude of an equilateral triangle is 18 inches. find the length of a side and the perimeter.
Answer by nerdybill(5404)

the altitude of an equilateral triangle is 18 inches. find the length of a side and the perimeter.
.
Let x = measure of one side
then from Pythagorean theorem:
18^2 +(x/2)^2 = x^2
18^2 +x^2/4 = x^2
18^2 = x^2 - x^2/4
18^2 = 3x^2/4
324 = 3x^2/4
108 = x^2/4
432 = x^2
20.785 inches = x
.
Perimeter is
3x = 3(20.785) = 62.354 inches